# Limitations of Data-Driven Spectral Reconstruction: Optics-Aware Analysis and Mitigation

### v3

@@ IV-D Discussion
The experiments conducted in this section clearly highlight several shortcomings with respect to the existing datasets: (1) they lack diversity in nuisance parameters such as noise and compression ratios, and (2) they lack scene diversity. While we show in this document the results for the largest available dataset (ARAD1K), the Supplement shows consistent results for all the datasets. Both of these aspects result in over-fitting and prevent the networks from learning the general spectral image restoration task. Next we specifically analyze the effect of metamerism; the analysis of the impact of optical aberrations will be deepened in Section VI.
@@ V-A Validation with metamers
We generate metamer datacubes (metamer data) from the original ARAD1K dataset (standard data) using the metameric black method [26]. For this set of experiments, we fix the coefficient $\alpha=0$ in Eq. (3). We choose a realistic imaging condition as used in Table LABEL:tab:unseen, Row 4, and keep it the same for both cases. The validation results on the ARAD1K dataset for existing pre-trained networks are summarized in Table V. We also visualize the reconstructed spectral images in five arbitrary bands (420 nm, 500 nm, 550 nm, 580 nm, and 660 nm) and spectra of two points in Fig. 4 for Scene ARAD_1K_0944 from the validation set.
@@
From the numerical results in Table V, it is apparent that all the existing methods experience catastrophic performance drop in terms of MRAE and SAM in the presence of metamers, which we call metameric failure. The MRAE (cf. Eq. (4)) may yield large values when big errors occur for dark ground-truth pixels (see the exemplary spectra in Fig. 4). The SAM values become large when the spectra are essentially dissimilar with each other. RMSE and PSNR do not capture the spectral differences as well, since they average out differences in the spatial and spectral dimensions.
@@ V-B Training with metamers
However, it is not sufficient to consider only a pair of standard and fixed metamer data. In reality, there are infinite metamers that project to the same color. As a second variant, we train the neural networks with random metamers generated on-the-fly as a spectral augmentation to enhance the spectral content of existing datasets. We vary the coefficient for the metameric black by setting $\alpha$ as a uniformly distributed random number in the range $[-1,2]$. During validation, we use both the standard validation data and their corresponding metamer data with fixed $\alpha=0$, which doubles the amount of the original validation data.
@@
As an example, we train MST++ and evaluate its validation performance over the training process. In Fig. 5, we show that it is no longer a good choice to use MRAE as the loss function [14] and evaluation metric [7], because it is completely overwhelmed by metamers. Instead, we find that L1 loss is a more stable loss function, so we train the network with L1 loss for three cases, no metamer (easy), fixed metamer (medium), and on-the-fly metamer (difficult). Nevertheless, we can see in Fig. 5 that the network fails in particular for the spectral accuracy SAM.
@@
We also train all other candidate networks with fixed and on-the-fly metamers. The results are summarized in Table VI. Again, the same performance drop applies to all networks. Finally, we show the results of the top-performing network, MST++ on the CAVE, ICVL, and KAUST datasets in Table VII. (See Supplemental Material for more results). As before, the performance drops similarly in the presence of metamers.
@@ V-C Discussion
The experiments conducted in this section clearly highlight the difficulties that the data-driven spectral recovery methods face with metamers: (1) lack of sufficient metameric data in current datasets, (2) training with metamers alone cannot mitigate the issue, and (3) spectral estimation from RGB data is indeed limited in the presence of metamers.
The limitations of spectral estimation from RGB data are ultimately not overly surprising – after all the projection from the high dimensional spectral space to RGB invariably destroys scene information that can be difficult to recover. Spatial context from underrepresented data does not contribute to the spectral estimation, because such information remains the same for metamers. However, our experiments show that this is indeed an issue faced by the state-of-the-art methods, which so far went unnoticed due to the under-representation of metamers in the datasets. This shortcoming will also affect other uses of the same datasets, for example in the training of reconstruction methods for spectral computational cameras [36, 9, 15]. Metameric adversary helps to identify this overlooked issue, whereas metameric augmentation has not yet found its way for contribution. It also underscores that, without side-channel information, no intrinsic property exists in the solely use of RGB images to distinguish between metamers, even when they are augmented. Note that this does not downplay the effects of metameric augmentation, since the problem formulation of reconstructing spectral information from RGB images in Eq. (1) is fundamentally limited. Once the problem is formulated in Eq. (2), metameric augmentation contributes to improving the network robustness. We will explore it further in the next section.
@@
VI Finding 3: The Aberration Advantage
As shown so far, the existing methods have difficulties distinguishing metamers in the ideal noiseless and aberration-free condition. In this section, we analyze what effect (if any) optical aberrations have on this situation, i.e., aberration-aware training [66]. To this end we train the networks in a realistic imaging condition with moderate noise level (npe = 1000), lossless PNG format, and aberrations from the same double Gauss lens as before [33]. In short, we simulate, through spectral ray tracing, the effect that an imperfect (i.e., aberrated) optical system has on the RGB image measured when observing a specific spectral scene. The details of this simulation can be found in the Supplemental Material.
In Fig. 6, we show an example with MST++ for the validation on SAM in two situations, one with fixed metamers, and the other with on-the-fly metamers. In each experiment, we compare the SAM difference with and without aberrations. As a reference, we also show the standard validation without metamers. As we can see, the realistic optical aberrations of the lens actually improve the spectral estimation in the presence of metamers as long as the aberrations are modeled in the training. With chromatic aberrations, the network can already distinguish fixed metamer pairs, achieving similar accuracy as the standard case. In the more aggressive case of on-the-fly metamers, chromatic aberrations also improve the spectral accuracy, compared with its no-aberration counterpart. To ablate on the effects of aberrations and metameric augmentation, we also plot the result with aberrations, but without metameric augmentation (purple lines). It implies that with aberrations only, spectral accuracy cannot be improved, because it simply means to reconstruct the hyperspectral images with more blurred RGB images. Again, this aberration advantage holds for all datasets (Table VIII). See Supplemental Material for details.
Refer to caption Refer to caption
Figure 6: Chromatic aberrations improve spectral accuracy. Left: fixed metamers. Right: on-the-fly metamers.
@@ VI Finding 3: The Aberration Advantage
VI-A Discussion
To understand why optical aberrations help improve the reconstruction, consider the simulated images in Fig. 7. The left and middle images are simulations of RGB images for metameric scene pairs, with the difference image on the right. The different spectra of the two scenes are affected differently by the optical aberrations, and therefore, although the scenes are metamers of each other, the RGB images are in fact different. In effect, the optical aberrations have encoded spectral information into the RGB image, which the networks can learn to distinguish, lending credibility to PSF engineering methods for hyperspectral encoding [36, 9, 15].
@@ VI-A Discussion
There are, however, also detrimental consequences of this effect, namely (1) “baked-in” aberrations in spectral datasets and (2) limited generalization to arbitrary RGB image sources. Regarding (1), it is worth noting that even expensive spectral cameras typically do not have diffraction-limited optics, i.e., their optical systems are aberrated. In order to avoid overfitting to these aberrations, future datasets should ideally be collected with a large variety of different spectral cameras. Regarding (2) we note that different RGB cameras produce very different aberrations, and often the aberrations also vary significantly across the image plane. Without further improvements in the reconstruction methods, this will result in a large device dependency and poor generalization of the methods.
VII Generalization analysis and effective spectral encoding
We have shown above three important findings in the RGB2HS formulation [7] through extensive experiments. The overfitting issues and metameric failure phenomenon contribute to the generalization difficulties in unseen data, particularly for metamers. The limitations could be mitigated with our proposed metameric augmentation along with aberration modeling as a spectral encoding scheme. Some other spectral encoding methods, such as CASSI, could also benefit from metameric augmentation. In Table IX, we show results on one of the best-performing network MST-L [13] trained the ARAD1K dataset and validated on the other three datasets. All the experiments are carried out with Poisson noise npe = 1000, and aberrations from the double Gauss lens (US20210263286A1). We adopt the CASSI settings in  [13].
As indicated by the results, RGB2HS cannot maintain its performance in PSNR and SAM for data in other datasets, even with metameric augmentation to account for metamers during training (see the RGB2HS columns). With the aid of aberrations and metameric augmentation, the reconstruction performance could be boosted for both standard and metamer data. In addition, evident from the CASSI results without metameric augmentation, CASSI offers overall better reconstruction quality, thanks to its spatial-spectral encoding design rooted in the compressive sensing theory. When metameric augmentation is applied to CASSI, its performance has also been boosted, proving the effectiveness of our metameric augmentation scheme. Note that the reconstruction quality varies among datasets, owing to the different spectral content in each dataset.
Effective spectral encoding is paramount to snapshot spectral imaging. The results of our extensive experiments reiterate the credibility to the computational camera approaches for spectral imaging, including both PSF engineering approaches [36, 9, 15] and compressed sensing methods like CASSI [62, 19]. However, learned reconstruction methods for these approaches also suffer from the same dataset issues as the methods analyzed in this paper, making the collection of large-scale, diverse spectral image data a matter of urgency. Our work reveals the underlying problems shared in most snapshot spectral imaging methods. Once they are well recognized, better solutions can be proposed to overcome these challenges in future works.
VIII Conclusion
@@
The major findings of our study are that (1) the limitations of current datasets lead to overfitting to both nuisance parameters (noise, compression), as well as limited scene content. (2) Metamerism in particular presents a challenge both in terms of under-representation in the datasets, and in terms of fundamental limits of spectral reconstruction from RGB input. (3) Metameric augmentation along with the targeted use of optical aberrations enables resolving the metamer issue.
Our results systematically demonstrate that it is impossible to accurately reconstruct spectrum solely from RGB images. In order to realize the dream of spectral estimation from arbitrary RGB sources, it is necessary to coherently and jointly diversify the spectral contents in hyperspectral image datasets, adopt side-channel information from the optical system, and embrace versatile spectral data augmentation methods to fully enable the power of networks in adaptation to whole families of aberrations.

### v4

@@ IV-D Discussion
The experiments conducted in this section clearly highlight several shortcomings with respect to the existing datasets. (1) They lack diversity in nuisance parameters such as noise and compression ratios. Our experiments show that when the RGB images are slightly modified by changing the nuisance parameters embedded in the image formation (e.g., noise level, compression factor, and optical aberrations), a significant drop in the performance is observed. (2) They lack scene diversity. Training modern deep neural networks via supervised learning typically demands large-scale datasets. However, our experiments reveal that reducing the training data volume leads to only a marginal or even no drop in spectral accuracy. This suggests that the dataset lacks diversity in its content. The cross-dataset validation experiments further show that each dataset has its own statistics. Neural networks trained on a single dataset exhibit limited generalization to other datasets, suggesting the lack of sufficient scene diversity in individual datasets. In summary, our experiments provide clear evidence that the performance degradation primarily stems from insufficient scene diversity in existing datasets. While we show here the results for the largest available dataset (ARAD1K), the Supplementary Material shows consistent results for all the datasets. Both of these aspects result in over-fitting and prevent the networks from learning the general spectral image restoration task. Next, we specifically analyze the effect of metamerism; the analysis of the impact of optical aberrations will be deepened in Section VI.
@@ V-A Validation with metamers
We generate metamer datacubes (metamer data) from the original ARAD1K dataset (standard data) using the metameric black method [35]. For this set of experiments, we fix the coefficient $\alpha=0$ in Eq. (3). We choose a realistic imaging condition as used in Table III, Row 4, and keep it the same for both cases. The validation results on the ARAD1K dataset for existing pre-trained networks are summarized in Table V. We also visualize the reconstructed spectral images in five arbitrary bands (420 nm, 500 nm, 550 nm, 580 nm, and 660 nm) and spectra of two points in Fig. 4 for Scene ARAD_1K_0944 from the validation set.
@@
HySAT std 0.326 0.047 29.4 0.127
met 61.516 0.095 26.2 0.594
HPRN std 0.524 0.104 22.0 0.130
met 33.439 0.102 22.8 0.574
SSTHyper std 0.314 0.058 27.7 0.117
met 48.427 0.089 25.8 0.575
MSFN std 0.328 0.055 28.3 0.119
met 51.846 0.090 26.0 0.573
GMSR std 0.484 0.075 24.9 0.138
met 53.063 0.109 23.6 0.547
SSRNet std 0.419 0.075 25.8 0.130
met 38.225 0.104 24.5 0.564
From the numerical results in Table V, it is apparent that all the existing methods experience catastrophic performance drop in terms of MRAE and SAM in the presence of metamers, which we call metameric failure. The MRAE (cf. Eq. (4)) may yield large values when large errors occur for dark ground-truth pixels (see the exemplary spectra in Fig. 4). The SAM values become large when the spectra are essentially dissimilar with each other. RMSE and PSNR do not capture the spectral differences as well, since they average out differences in the spatial and spectral dimensions.
@@ V-B Training with metamers
However, it is not sufficient to consider only a pair of standard and fixed metamer data. In reality, there are infinite metamers that project to the same color. As a second variant, we train the neural networks with random metamers generated on-the-fly as a spectral augmentation to enhance the spectral content of existing datasets. We vary the coefficient for the metameric black by setting $\alpha$ as a uniformly distributed random number in the range [1,2][-1,2]. During validation, we use both the standard validation data and their corresponding metamer data with fixed $\alpha=0$, which doubles the amount of the original validation data.
@@
As an example, we train MST++ and evaluate its validation performance over the training process. In Fig. 5, we show that it is no longer a good choice to use MRAE as the loss function [15] and evaluation metric [7], because it is completely overwhelmed by metamers. Instead, we find that L1 loss is a more stable loss function.
$\mathcal{L}_{1}\left(\hat{{\mathbf{X}}},{\mathbf{X}}\right)=\frac{1}{KMN}\sum_{k,i,j}\left|\hat{{\mathbf{X}}}_{k,i,j}-{\mathbf{X}}_{k,i,j}\right|.$ (8)
We then train the network with L1 loss for three cases, no metamer (easy), fixed metamer (medium), and on-the-fly metamer (difficult). Nevertheless, we can see in Fig. 5 that the network fails in particular for the spectral accuracy SAM.
@@
HySAT no 0.295 0.036 32.6 0.078
fixed 7.450 0.063 29.9 0.270
on-the-fly 12.806 0.087 26.5 0.283
HPRN no 0.249 0.038 32.7 0.076
fixed 5.633 0.060 31.6 0.221
on-the-fly 12.197 0.080 26.7 0.263
SSTHyper no 0.261 0.036 32.4 0.083
fixed 10.070 0.060 30.5 0.285
on-the-fly 12.194 0.083 27.3 0.276
MSFN no 0.283 0.043 30.4 0.078
fixed 8.609 0.065 29.1 0.286
on-the-fly 12.198 0.096 24.9 0.288
GMSR no 0.333 0.049 29.0 0.114
fixed 10.612 0.065 29.4 0.302
on-the-fly 11.674 0.096 24.9 0.300
SSRNet no 0.354 0.052 29.0 0.098
fixed 6.678 0.072 28.9 0.286
on-the-fly 13.359 0.102 25.0 0.291
We also train all other candidate networks with fixed and on-the-fly metamers. The results are summarized in Table VI. Again, the same performance drop applies to all networks. Finally, we show the results of the top-performing network, MST++ on the CAVE, ICVL, and KAUST datasets in Table VII. (See Supplementary Material for more results). As before, the performance drops similarly in the presence of metamers.
@@ V-C Discussion
The experiments conducted in this section clearly highlight the difficulties that the data-driven spectral recovery methods face with metamers: (1) lack of sufficient metameric data in current datasets, (2) training with metamers alone cannot mitigate the issue when the problem is formulated by Eq. (1), and (3) spectral estimation from RGB data is indeed limited in the presence of metamers.
@@
The limitations of spectral estimation from RGB data are ultimately not overly surprising – after all the projection from the high dimensional spectral space to RGB invariably destroys scene information that can be difficult to recover. Spatial context from underrepresented data does not contribute to the spectral estimation, because such information remains the same for metamers. However, our experiments show that this is indeed an issue faced by the state-of-the-art methods, which so far went unnoticed due to the under-representation of metamers in the datasets. This shortcoming will also affect other uses of the same datasets, for example in the training of reconstruction methods for spectral computational cameras [47, 10, 20]. A metameric adversary helps to identify this overlooked issue and avoid unrealistically high numerical scores for existing systems. It also underscores that, without side-channel information, no intrinsic property exists in RGB images to distinguish between metamers, even when they are augmented. Note that this does not downplay the effects of metameric augmentation, since the problem formulation of reconstructing spectral information from RGB images in Eq. (1) is fundamentally limited. Once the problem is formulated in Eq. (2), metameric augmentation contributes to improving the network robustness. We will explore it further in the next section.
VI Finding 3: The Aberration Advantage and Effective Spectral Encoding
VI-A Aberration-Aware Training with Metameric Augmentation
As shown so far, the existing methods have difficulties distinguishing metamers in the ideal noiseless and aberration-free condition. In this section, we analyze what effect (if any) optical aberrations have on this situation, i.e., aberration-aware training [94]. To this end, we train the networks in a realistic imaging condition with moderate noise level (npe = 1000), lossless PNG format, and aberrations from the same double Gauss lens as before [44]. In short, we simulate, through spectral ray tracing, the effect that an imperfect (i.e., aberrated) optical system has on the RGB image measured when observing a specific spectral scene. The details of this simulation can be found in the Supplementary Material.
In Fig. 6, we show an example with MST++ for the validation on SAM in two situations, one with fixed metamers, and the other with on-the-fly metamers. As a reference, we also show the standard validation without metamers as done in previous works (thin dashed black lines). As we can see, the realistic optical aberrations of the lens actually improve the spectral estimation in the presence of metamers as long as metamers are modeled in the training. With chromatic aberrations combined with metameric augmentation, the network can already distinguish fixed metamer pairs, achieving similar accuracy as the standard case. In the more aggressive case of on-the-fly metamers, chromatic aberrations also improve the spectral accuracy, compared with their no-aberration counterparts. Again, this aberration advantage holds for all datasets (Table VIII). See Supplementary Material for details.
@@ VI-A Aberration-Aware Training with Metameric Augmentation
Refer to caption Refer to caption
Figure 6: Chromatic aberrations improve spectral accuracy. Left: fixed metamers. Right: on-the-fly metamers. aberr.: aberrations; met-aug.: metameric augmentation.
We carry out an ablation study to examine all possible combinations of aberrations and metameric augmentation, as shown in Fig. 6. This leads to 4 situations: (1) Training without aberrations, and without metameric-augmentation (purple lines). It is the existing training method that fails to combat metamers, as we have demonstrated in Section V. (2) Training with aberrations, but without metameric augmentation (green lines). It corresponds to simply a more blurry RGB image without considering metamers, which fails similarly. (3) Training without aberrations, but with metameric augmentation (blue lines). It slightly improves the spectral reconstruction, owing to the fact that noise may introduce slight differences in the RGB images. (4) Training with both aberrations and metameric augmentation. With the combination of aberrations and metameric augmentation, the network can learn the differences in RGB images between metamers, achieving similar performance as existing methods where metamers are actually not evaluated. These results clearly demonstrate again that the spectral reconstruction solely from RGB images fails for metamer data. The problem can be more effectively formulated only when spectral information, such as optical aberrations, is taken into account.
@@
To understand why optical aberrations help improve the reconstruction, consider the simulated images in Fig. 7. The left and middle images are simulations of RGB images for metameric scene pairs, with the difference image on the right. The different spectra of the two scenes are affected differently by the optical aberrations, and therefore, although the scenes are metamers of each other, the RGB images are in fact different. With optical aberrations, spectral information spreads out into adjacent pixels spatially at the cost of slightly making the RGB images blurry. The networks then see these color differences from metamers during training, and are able to correctly learn the mapping from metameric spectra to colors. We visualize the reconstructed spectra for the two example points in Fig. 4 with the MST++ model trained with both aberrations and on-the-fly metameric augmentation in Fig. 8. Now it is clear that when the network is trained with aberrations and metameric augmentation, it can tell the metameric spectra apart, which is not possible otherwise. In effect, the optical aberrations have encoded spectral information into the RGB image, which the networks can learn to distinguish, lending credibility to PSF engineering methods for hyperspectral encoding [47, 10, 20]. We also note that incidental aberrations of a lens are only a weak spectral encoding. Better performance would require optimizing the spectral encodings deliberately.
Figure 8: Reconstructed metameric spectra with MST++ trained with aberrations and metameric augmentation. The two example spectra are exactly the same as in Fig. 4.
VI-B Effective Spectral Encoding
The above experiments point to a better formulation of the spectral reconstruction problem where spectral encoding plays a key role. Optical aberrations, however, are usually minimized in camera lenses, making them imperfect candidates for effective spectral encoding. Existing works have explored some deliberate use of dispersive optical elements for this purpose, including diffraction rotation [47] and grating [10]. However, their performance has historically not been analyzed for metamers either. To this end, we test such spectral encoding schemes on a real challenging scene as shown earlier in Fig. 1. In Fig. 9, we compare four spectral encoding conditions: None (no encoding), Diffraction Rotation (used in [47]), Double Gauss aberrations (used above), and Grating (used in [10]). We train MST++ with metameric augmentation for these spectral encodings, and the SAM results in Fig. 9(b) show that spectral encodings indeed improve the overall spectral accuracy compared with no spectral encoding. The corresponding spectral PSFs are shown in Fig. 9(c). The reconstructed spectra for the same two points in Fig. 1 are shown in Fig. 9(d). Without spectral encoding, the spectral accuracy diverges, while all the spectral encodings improve the spectral quality. Interestingly, different spectral encoding schemes lead to varying spectral accuracy. Diffraction rotation tends to separate the metameric colors more, but the overall SAM is worse than aberrations and grating. We highlight that such challenging metameric spectra have not been properly evaluated in previous works. Although these spectral encoding schemes have been proposed, they are not optimized to deal with metamers yet. Again, the primary culprits are the dataset limitations we have pointed out – metamers are highly underrepresented in existing datasets. While the CAVE dataset includes such examples, they are present in only limited quantities. To achieve better spectral reconstruction in such challenging situations, an effective spectral encoding and a powerful neural network should be trained on a large-scale and diverse dataset in which metamers are well represented. This constitutes a critical yet unresolved challenge in the field of data-driven spectral reconstruction.
Figure 9: Comparison of different spectral encoding methods. (a) The fake_and_real_food_ms scene from CAVE. (b) Training performance in SAM for different encoding methods. (c) PSFs for different spectral encodings: none, diffraction rotation [47], double Gauss aberration, and grating [10]. (d) Reconstructed spectra for the two points denoted in (a).
@@ VI-B Effective Spectral Encoding
Other spectral encoding methods, such as CASSI, could also benefit from metameric augmentation. In Table IX, we show results on one of the best-performing networks, MST-L [14], trained on the ARAD1K dataset and validated on the other three datasets. All the experiments are carried out with Poisson noise npe = 1000, and aberrations from the double Gauss lens (US20210263286A1). We adopt the CASSI settings in  [14].
As indicated by the results, RGB2HS cannot maintain its performance in PSNR and SAM for data from other datasets, even with metameric augmentation to account for metamers during training (see the RGB2HS columns). With the aid of aberrations and metameric augmentation, the reconstruction performance could be boosted for both standard and metamer data. In addition, evident from the CASSI results without metameric augmentation, CASSI offers overall better reconstruction quality, thanks to its spatial-spectral encoding design rooted in the compressive sensing theory. When metameric augmentation is applied to CASSI, its performance has also been boosted, proving the effectiveness of our metameric augmentation scheme. Note that the reconstruction quality varies among datasets, owing to the different spectral content in each dataset.
VI-C Discussion
We have demonstrated through aberration-aware training with metameric augmentation that only when optical aberrations are considered in the image formation, the problem of spectral reconstruction from RGB images can be better formulated. Our ablation experiments further prove that it is not sufficient to only model metamers without employing spectral encodings. In addition, we highlight that optical aberrations are incidental but not deliberate spectral encodings, so their effect is still limited. We further examine alternative spectral encoding methods, such as PSF engineering approaches [47, 10, 20], and compressed sensing methods like CASSI [79, 25]. The results of our extensive experiments reiterate the credibility to such computational camera approaches. However, learned reconstruction methods for these approaches also suffer from the same dataset issues as the methods analyzed in this paper, making the collection of large-scale, diverse spectral image data a matter of urgency. These new datasets in turn will enable the design of improved optical encodings in computational spectral imaging systems without overfitting to specific scenarios.
VII Conclusion
@@
The major findings of our study reveal important yet previously overlooked limitations in this research direction. (1) The limitations of current datasets lead to overfitting to both nuisance parameters (noise, compression), as well as limited scene content. (2) Metamerism in particular presents a challenge both in terms of under-representation in the datasets, and in terms of fundamental limitations of spectral reconstruction from RGB input. (3) Metameric augmentation along with the targeted use of optical aberrations paves the way to combating the metamer issue, though more effective spectral encodings are demanded to solve the challenge.
Our results systematically demonstrate that it is impossible to accurately reconstruct spectra solely from RGB images. In order to realize the dream of spectral estimation from arbitrary RGB sources, it is necessary to coherently and jointly diversify the spectral contents in hyperspectral image datasets, adopt side-channel information from the optical system, and embrace versatile spectral data augmentation methods to fully enable the power of networks in adaptation to whole families of spectral encodings. We argue that addressing these foundational issues is imperative. Continuing to propose new network designs without rethinking the misdefined problem formulation will fall into the same fundamental shortcomings. Clarifying these limitations will enable the community to focus on solving the real challenges in snapshot spectral imaging.
The dataset limitations we point out in this work may also apply to other spectral reconstruction problems using the referenced datasets, such as CASSI, PSF engineering, and multispectral-hyperspectral fusion. In particular, the same metamerism issue has not yet been extensively evaluated in such domains either. Our findings underscore the broader importance of effective spectral encodings in such snapshot spectral imaging problems. The proposed metameric augmentation technique could inform future directions in optical design, network design, and, more importantly, their joint optimization to cope with metamers.
Acknowledgments
This work was supported by the KAUST Individual Baseline Funding and Center of Excellence for Generative AI. Seung-Hwan Baek was supported by the National Research Foundation of Korean (NRF) grant funded by the Korea Goverment (MSIT) (RS-2024-00438532, RS-2023-00211658) and Basic Science Research Program through the NRF funded by the Ministry of Education (2022R1A6A1A03052954).